OpenStudy (anonymous):

. Archie's gas tank is 1/3 full. If Archie adds 3 gallons of gas to the tank, it will be 1/2 full. What is the capacity in gallons of Archie's tank?

OpenStudy (anonymous):

lets say x gallons is the capacity of tank.... initially, tank has x/3 gallons after Archie adds 3 more gallons, x/3+3=x/2 solve for x

OpenStudy (nurali):

\[\frac{ 1 }{ 3 }x+3=\frac{ 1 }{ 2 }x\] Multiply each term by 6 2x + 6(3) = 3x 18 = 3x - 2x x = 18